INFRARED CAVITY RINGDOWN SPECTROSCOPY OF JET-COOLED PAHS: A COMPARISON WITH MATRIX SPECTRA

\underline{A. J. HUNEYCUTT}, R. N. CASAES, B. J. McCall, R. J. SAYKALLY, Department of Chemistry, University of California, Berkeley, CA 94720; C.-Y. CHUNG, and Y.-P. LEE, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan..

Infrared absorption spectra of the CH stretching region were observed for naphthalene, anthracene, phenanthrene, pyrene, and perylene using a heated supersonic slit source and cavity ringdown spectroscopy. Results are compared closely with 10 K Ar matrix spectra to determine general matrix perturbation effects for this class of molecules. Fundamental transitions in the matrix spectra were subject to spectral shifts of up to 3.0 cm^{−1} and band widths were generally broader than the jet-cooled spectra by up to 80%. Weak features not predicted by theory were observed in both Ar matrix and gas-phase spectra with similar relative intensities which suggest assignment to overtones and combination bands.